OpenStudy (anonymous):

A factory worker does decorative stitching on jeans. He stitched 37 pairs of jeans from 8:00 a.m. to 12:00 noon. If he took a 45-minute break during that time, which equation can be used to determine j, the average number of pairs of jeans he completed per hour of work?

OpenStudy (anonymous):

a. j = 37 ÷ (4 − 0.75) b. j = (37 − 0.75) × 4 c. j = (37 − 0.75) ÷ 4 d. j = 37 × (4 + 0.75) e. j = 37 × (4 − 0.75)

OpenStudy (texaschic101):

he stitched 37 pair of jeans in 4 hrs he took 45 minute break....which is 75% of an hr, which is 0.75 so : j = 37 divided by (4 - 0.75)

OpenStudy (anonymous):

Thank you for explaining what the .75 meant i was kind of confused about that

OpenStudy (anonymous):

The bill from your plumber was $150. The cost for labor was $40 per hour. The cost for materials was $50. How many hours did the plumber work? Answer in units of hours